# leftmargin=0.15cm, rightmargin=0.15cm, font=italic, vskip=0.10cm, indentfirst=false

### v4

@@ Abstract
Software signing is the most robust method for ensuring the integrity and authenticity of components in a software supply chain. However, traditional signing tools place key management and signer identification burdens on practitioners, leading to both security vulnerabilities and usability challenges. Next-generation signing tools such as Sigstore have automated some of these concerns, but little is known about their usability and adoption dynamics. This knowledge gap hampers the integration of signing into the software engineering process.
@@
To fill this gap, we conducted a usability study of Sigstore, a pioneering and widely adopted exemplar in this space. Through 18 interviews, we explored (1) the factors practitioners consider when selecting a signing tool, (2) the problems and advantages associated with the tooling choices of practitioners, and (3) practitioners’ signing-tool usage has evolved over time. Our findings illuminate the usability factors of next-generation signing tools and yield recommendations for toolmakers, including: (1) enhance integration flexibility through officially supported plugins and APIs, and (2) balance transparency with privacy by offering configurable logging options for enterprise use.
I Introduction
The reuse of software components in modern products creates complex supply chains and expands the attack surface for these systems [1, 2, 3]. A core strategy for mitigating this risk is to establish provenance—verifying actor authenticity and artifact integrity—through software signing [4, 5]. Software signing is the strongest guarantee of provenance [6] and is widely advocated by academia, industry consortia, and government regulations [7, 8, 9]. High-profile supply-chain attacks—such as the SolarWinds breach—have further underscored the need for stronger, more automated signing infrastructure. In response, tools have evolved into two classes: traditional solutions (e.g., OpenPGP [10]), which delegate key management and signer identification to practitioners, and next-generation, identity-based tools (e.g., Sigstore [11]), which further automate the process via short-lived keys and external identity providers.
Tooling has significantly advanced various technological fields [12, 13, 14, 15] and has often played a key role in the adoption of these technologies [16, 17]. There is a continual need to evaluate the adoption practices and usability of such tools to inform practitioners, potential users, researchers, and toolmakers of their effectiveness in different contexts [18]. Traditional software signing tools have been studied in this way, most notably in several “Why Johnny Can’t Encrypt” inspired studies[19, 20, 21]. Although next-generation signing tools promise greater automation, there remains a lack of comprehensive usability and adoption analyses for these solutions. Such studies could assess how automation affects adoption suitability, elicit practitioners’ desired improvements, and guide future tool design and research.
Our work offers the first examination of next-generation software signing tools’ usability as a critical factor influencing their adoption and practice. Building on and improving previous methodologies — which relied on novice users and summative usability assessments criticized for their lack of actionable feedback[18, 22] — we conducted interviews with 18 experienced security practitioners to study current software signing tools, focusing on Sigstore as a case study. We analyzed our data with a formative usability framework [23], focusing on two usability dimensions: practitioners’ experiences with Sigstore and its usability relative to other signing tools. From these interviews, we distilled three key themes: the evolution and motivations behind tool adoption, the limitations and challenges of currently adopted tools, and the specific implementation practices shaping real-world signing workflows.
Our results highlight the usability factors practitioners consider when adopting next-generation software signing tools. This serves as formative feedback for tool designers and researchers, offering insights into the drivers of tool usage. We present these findings through a clear articulation of practitioners’ challenges and the merits of using particular tools. Additionally, we found that practitioners’ decisions to adopt a software signing tool are influenced by less commonly discussed factors, such as the tool’s community, relevant regulations and standards, and certain functionalities of next-generation Software signing tools.
@@ I Introduction
Significance of Study: Our work targets Sigstore—a pioneer next-generation signing tool that has seen rapid uptake across major open-source ecosystems[24, 25, 26], cloud-native projects, and corporations[27, 28, 29, 30, 31], involving a large number of engineers worldwide. By unpacking the usability factors that drive or hinder Sigstore’s adoption, we not only inform the ongoing refinement of its automation workflows (e.g., certificate issuance and key management) but also provide a template for evaluating and improving similar Next-generation(identity-based) signing solutions. Also, since Sigstore exemplifies a broader class of tools that embed security into automated build and deployment processes, our findings can be potentially generalized to guide the design and implementation of other automated software-supply-chain technologies, ultimately advancing the usability—and thus the security—of automated software engineering practices.
II Background & Related Works
In this section, we review software signing technologies (section II-A) and usability assessments (section II-B).
II-A Software Signing
Software signing ensures the authorship and integrity of a software artifact by linking a maintainer’s cryptographic signature to their software artifact. The implementation of software signing is shaped by several factors: the software artifacts to be signed (e.g., driver signing[32]), the root-of-trust design (e.g., public key infrastructure/PKI [33]), and ecosystem policies (e.g., PyPi[34] and Maven[35]).
II-A1 Software Signing Tools
@@
Next-generation (identity-based) signing tools automate key management by issuing short-lived, ephemeral certificates tied to a signer’s identity via external OpenID Connect (OIDC) or OAuth providers. This “keyless signing” approach — exemplified by Sigstore [11], OpenPubKey [37], and SignServer [38] — eliminates long-term key handling and embeds transparency through append-only logs, improving usability and security. We discuss the architecture and workflow for next-generation signing tools in section IV-B using the case of Sigstore.
II-A2 Effects of Signing Tools on Software Signing Adoption
@@
Previous studies recognize that better tooling influences software-signing adoption. For example, Schorlemmer et al. [6] performed a measurement study of several software package registries, identifying tooling as a key determinant of signature quality. Similarly, Kalu et al. [39] explore implementation and adoption frameworks in commercial settings, highlighting tool usability, compatibility, and integration challenges. While these works emphasize broad usability factors, they offer little detailed guidance on what makes a signing tool usable, overlooking the fine-grained technical and workflow requirements that drive real-world tool adoption.
II-B Usability Studies in Software Signing
II-B1 Usability – Definition, Theory & Evaluation Frameworks
@@
Tools enable engineers to adopt new software development practices [40]. A usable tool is one whose functionality effectively supports its intended purpose—achieving product goals [18, 41]. Studies assessing implementation strategies of tools are typically framed as usability evaluations [42].
@@
To evaluate a tool’s usability, studies typically follow a usability evaluation framework [42]. There are two kinds: summative, which measures a tool’s effectiveness [43, 44], and formative, which provides iterative feedback for improvement [18, 45]. Given the aims of this work, we took a formative approach. We considered several formative frameworks [46, 47, 48, 49] and found that Cresswell et al.’s four-factor framework [23] helped explain our data the best. That framework is summarized in Figure 1.
Figure 1: Cresswell’s Usability Framework [23]. It assesses usability based on four factors. Technological (T) concerns the technical characteristics of the system (e.g., performance, ease of use) and data management (availability, integrity). Social/Human (P) relates to users and their interactions with the tool, e.g., user satisfaction, engagement, prior experience, and correct use of features. Organizational Context (O) are factors internal to participants’ organization, e.g., training and leadership support. Macroenvironmental (M) are external influences, e.g., regulations, economics, and vendor relations.
II-B2 Empirical Studies
@@
The usability of traditional signing tools—often studied in the context of encryption—remains a significant challenge. Whitten and Tygar’s seminal study, “Why Johnny Can’t Encrypt” [19], found PGP 5.0’s interface too complex for non-experts, a problem that persisted in PGP 9 [20], especially around key verification, transparency, and signing. Subsequent work [50, 51, 52] has underscored the public-key model’s inherent complexity. Research comparing manual versus automated signing tools has yielded mixed results: some works observed no clear usability advantage [53, 54], whereas Atwater et al. showed a user preference for automated PGP-based solutions [55]. Although informative, these studies focus on traditional encryption scenarios (e.g., email and messaging), limiting their applicability to broader software-signing workflows.
These studies exhibit weaknesses in their scope, methodology, and the relevance of their findings to software signing. First, they primarily focus on evaluating the usability of signing tooling in the context of signing communication (e.g., email, messages), limiting their application to signing software (i.e., securing the software supply chain). Furthermore, much of this research is outdated, reflecting decades-old technologies and failing to incorporate contemporary practitioner experiences. Their methodology predominantly employs summative usability assessments that rely on novice volunteers, which limits the applicability of their findings for guiding tool enhancements [56, 57].
III Knowledge Gaps & Research Questions
Existing literature on software signing (see section II) has examined practice and adoption across private and open-source ecosystems, highlighting the importance of tool choice in signature quality. Prior usability studies have focused exclusively on traditional signing tools and have been limited to evaluating user-interface design. However, next-generation signing tools remain understudied, creating a need for:
  • A holistic evaluation of how software signing tools influences software-signing adoption and practice, and how usability affects this process.
  • Practitioner perspectives on the nuanced role of tooling in shaping software-signing use and practice. Current research (see section II-B2) is limited because it primarily relies on experiments with novice users, while studies in section II-A2 use quasi-experiments or general practitioner interviews without a specific focus on signing tools.
In light of these gaps, a comprehensive usability evaluation of next-generation signing tools would offer valuable insights into their role in adoption, improve their design, and enhance their adaptability.
Research Questions: Building on identified gaps in next-generation tooling usability and its impact on adoption, we frame our research around Sigstore—our chosen next-generation signing tool case study—and ask:
@@ III Knowledge Gaps & Research Questions
    What strengths and weaknesses of Sigstore influence its adoption in practice?
@@
    What factors influence practitioners’ decisions to change the software signing tools they use?
IV Methodology
We selected a case study method to answer our RQs. Our methodology is illustrated in Figure 2. This section provides our design rationale (section IV-A), defines the case study context (section IV-B), discusses instrument design (section IV-C), data collection (section IV-D), and analysis (section IV-E).
@@ IV Methodology
IV-A Rationale for Case Study
The case study is a research method used to investigate a phenomenon within its real-life context  [58, 59, 60, 61]. We selected the case study method for our research because it allows for an in-depth exploratory examination of our chosen case, enabling a deeper understanding of the usability factors that influence the adoption of a software signing tool in practice, rather than a broad but superficial cross-analysis [62]. In selecting Sigstore, we adhered to the criteria outlined by Yin [59] and Crowe et al. [60]. Specifically, we considered the intrinsic value and uniqueness of the case study unit (high), access to data (adequate), and risks of participation (low). In section VI we discuss how our findings may generalize.
Figure 2: Study Methodology. We developed our interview protocol by reviewing Sigstore grey literature and the academic literature. Finally, we analyzed our results using a thematic analysis informed by Cresswell’s formative usability evaluation framework (Figure 1).
IV-B Case Study Context – Sigstore
@@
Use in Practice: Sigstore launched in 2021. It provides keyless, ephemeral keys and signer authentication mechanisms to simplify the signing process used by traditional tools. Within 13 months of its public release, Sigstore recorded 46 million artifact signatures [25]. It has significant adoption and support from open-source ecosystems (e.g., PYPI [26], Maven [24], NPM [63]) and many companies (e.g., Autodesk [31], Verizon [29], Yahoo[64]). Thus, Sigstore is a major next-generation software signing platform, making it a suitable context for examining usability in modern signing tools.
Components & Workflow: We detail Sigstore’s architecture and signing workflow, to familiarize the reader with this example of next-generation software signing tools, and to interpret participants’ statements referring to these components.
Sigstore’s primary components are [11]:
  • Cosign: CLI and Go library to create, verify, and bundle signatures and attestations.
  • OIDC Identity Provider: OpenID Connect services (e.g., GitHub, Google, Microsoft) that authenticate signers and issue short-lived identity tokens.
@@
  • Monitors and Verifiers: Services or tools that validate the signer’s identity, certificate status, and signature integrity.
Figure 3 shows the Sigstore software signing workflow. First, the signer authenticates with an OIDC provider to obtain a short-lived token. Next, Fulcio verifies this token and issues a one-time signing certificate. The signer then uses Cosign to apply that certificate to the software artifact. Finally, Rekor logs the certificate and signature, and verifiers retrieve entries from both Fulcio’s log and Rekor to confirm the signer’s identity and ensure the certificate and signature remain valid.
Figure 3: Sigstore Signing Workflow. The software author requests a certificate from a certificate authority, which confirms the signer’s identity through an identity provider. The signature and signer’s identity are recorded in a transparency log, which the verifier can monitor to confirm the validity of the signature upon downloading the signed package. See section -K for legible figure
Commonality with other Next-Gen Tools: Sigstore, like other next-generation signing tools such as OpenPubKey [37], AWS Signer [65], and SignServer [38], shares core functionality in signer identity management and automated key handling. However, unlike Sigstore, most of these tools do not include a built-in certificate authority (e.g., OpenPubKey) or a transparency log (e.g., OpenPubKey, SignServer). Instead, such services must be added via external integrations. While our focus on Sigstore allows for an in-depth case analysis, many of its design principles—such as identity-based authentication and ephemeral keys—are shared by its peers, suggesting potential generalizability across next-generation signing tools
IV-C Instrument Design & Development
Following guidance from Saldaña [66], we used semi-structured interviews to examine various aspects of tooling usability, e.g., implementation challenges, perceived benefits, and comparison to other signing tools. This approach let us pose consistent questions across all subjects, with the ability to probe about unique aspects of each subject’s circumstances.
To develop our interview instrument, we did not bind our questions to any prior usability framework. Sigstore is a nascent technology and in such contexts an exploratory design is recommended [59, 58]. Only after this exploratory phase did we map our findings onto a formal usability framework, ensuring it reflected practitioners’ lived experiences rather than constraining data collection to preconceived categories. Therefore, to construct the interview protocol, we:

### v5

@@ Abstract
Software signing is the most robust method for ensuring the integrity and authenticity of components in a software supply chain. Traditional signing tools burdened practitioners with key management and signer identification, creating both usability challenges and security risks. A new class of next-generation signing tools have automated many of these concerns, but little is known about their usability and its effect on adoption and effectiveness in practice. A usability evaluation can clarify the extent to which next-generation designs succeed and highlight priorities for improvement.
@@
To fill this gap, we conducted a usability study of Sigstore, a pioneering and widely adopted exemplar of next-generation signing. Through interviews with 17 industry experts, we examined (1) the problems and advantages associated with practitioners’ tooling choices, (2) how and why their signing-tool usage has evolved over time, and (3) the contexts that cause usability concerns. Our findings illuminate the usability factors of next-generation signing tools and yield recommendations for toolmakers, adopting organizations, and the research community. Notably, components of next-generation tooling exhibit different levels of maturity and readiness for adoption, and integration flexibility is a common pain point but potentially mitigable through plugins and APIs. Our results will help next-generation signing toolmakers further strengthen software supply chain security.
1 Introduction
The reuse of software components in modern software production creates complex supply chains that are susceptible to the unauthorized introduction of code [100, 84, 11, 48]. To mitigate this risk, engineers need provenance—evidence of actor authenticity and artifact integrity [15]. Software signing provides the strongest guarantee of provenance currently available [70] and has thus been widely advocated by academia, industry consortia, and government regulators [92, 22, 21, 19].
Like other security practices, the effectiveness of software signing in practice depends on the availability and usability of tools that support it. Traditional signing tools such as OpenPGP are famously challenging to use [98], with several “Why Johnny Can’t Encrypt”–inspired studies consistently revealing usability challenges that limit adoption [99, 67, 66]. In recent years, next-generation, identity-based systems (e.g., Sigstore [104]) have been introduced that automate much of the process using short-lived keys and external identity providers [26]. However, despite growing adoption driven by recent supply chain attacks and efforts to lower usability barriers [104, 69, 55], we lack evidence about the usability of these next-generation tools and need to understand how that affects adoption in practice [51]. The innovations in next-generation approaches to signing creates a pressing need to evaluate how usability concerns affect adoption and effectiveness. A usability evaluation can clarify the extent to which next-generation designs succeed, highlight how usability compares to traditional approaches, and surface priorities for improvement.
To this end, we offer the first empirical study of the usability of next-generation software signing tools. Our work examines Sigstore—a pioneer next-generation signing tool that has seen rapid uptake across major open-source ecosystems[81, 36, 49], cloud-native projects, and corporations[96, 45, 76, 75, 77], involving a large number of engineers worldwide. We interviewed 17 experienced security practitioners to investigate the usability concerns that drive or discourage Sigstore’s adoption. We adopted an exploratory data analysis approach, because this study addresses a new class of tools that has not previously been analyzed. Our analysis centered on two aspects: practitioners’ direct experiences with Sigstore and their perceptions of its usability relative to other signing tools.
Our study highlights the usability factors practitioners weigh when adopting next-generation software signing tools. We show that while next-generation signing tools ease traditional challenges, adoption is still constrained by integration hurdles, privacy concerns, and organizational factors. Second, our results provide adopting organizations with insight into the readiness levels of these tools, helping them manage adoption risks. Finally, we offer toolmakers and designers concrete directions for improving next-generation signing tools. Together, these findings provide formative feedback situated in real-world usage contexts.
@@ 1 Introduction
  3. 3.
    We discuss how organizational deployment contexts influence the usability concerns raised by practitioners.
Significance: Our work contributes to the security of software signing tools and, more broadly, to software supply chains. By unpacking the usability factors that promote and hinder Sigstore’s adoption, we inform the ongoing refinement of its workflows and provide a template for evaluating and improving similar next-generation (identity-based) signing solutions. Because Sigstore exemplifies other software signing tools, our findings can be generalized to guide the design and implementation of this class of software, ultimately advancing usability (and thus security) across the software supply chain.
2 Background & Related Works
In this section, we review software signing technologies (section˜2.1) and usability assessments (section˜2.2).
2.1 Software Signing
Software signing ensures the authorship and integrity of a software artifact by linking a maintainer’s cryptographic signature to their software artifact. From literature, the implementation of software signing is shaped by several factors: the software artifacts to be signed (e.g., driver signing[24]), the root-of-trust design (e.g., public key infrastructure/PKI [105]), and ecosystem policies (e.g., PyPi[87] and Maven[80]).
Software Signing Tools
@@ 2.1 Software Signing
Next-generation (identity-based) signing tools automate key management by issuing short-lived, ephemeral certificates tied to a signer’s identity via external OpenID Connect (OIDC) or OAuth providers. This “keyless signing” approach, exemplified by Sigstore [104], OpenPubKey [26], and SignServer [73], eliminates long-term key handling and embeds transparency and auditability through append-only logs. We discuss the architecture and workflow for next-generation signing tools in section˜4.2 using the case of Sigstore.
Effects of Signing Tools on Software Signing Adoption
@@
Previous studies[55, 70, 42] recognize that better tooling influences software-signing adoption. For example, Schorlemmer et al. [70] performed a measurement study of several software package registries, identifying tooling as a key determinant of signature quality. Similarly, Kalu et al. [42] explore implementation and adoption frameworks in commercial settings, highlighting tool usability, compatibility, and integration challenges. While these works emphasize broad usability factors, they offer little detailed guidance on what makes a signing tool usable, overlooking the fine-grained technical and workflow requirements that drive real-world tool adoption.
2.2 Usability Studies in Software Signing
2.2.1 Usability — Definition and Evaluation Approaches
@@
Tools enable engineers to adopt new software development practices [85]. A usable tool is one whose functionality effectively supports its intended purpose, achieving product goals [51, 1]. Studies assessing implementation strategies of tools are typically framed as usability evaluations [4].
@@
To evaluate a tool’s usability, studies typically follow a usability evaluation framework [4]. There are two kinds of approaches. Summative work measures a tool’s effectiveness [39, 16], e.g., by task success rates, completion time, and user satisfaction. Formative assessments provide feedback to guide ongoing improvement [51, 93]. Given our aims, we took a formative approach. We considered several formative frameworks [86, 95, 89, 47] and found that Cresswell et al.’s four-factor framework [41] best explained our data. That framework is summarized in Figure˜1.
Usability depends on the context in which a tool is used. Many factors, such as organizational objectives, personnel, policies, and compliance requirements, can make a tool usable in one setting and impractical in another [99, 50]. This contextual aspect informs our study design.
Figure 1: Study Methodology and Usability Framework. We developed our interview protocol by reviewing academic and grey literature on software signing, then conducted 17 practitioner interviews. Data were analyzed through thematic analysis and mapped onto Cresswell’s formative usability evaluation framework [41]. The framework organizes usability concerns into four factors: Technological (T) — technical characteristics such as performance, ease of use, and data integrity; Social/Human (P) — user experience, and correct use of features; Organizational (O) — internal factors such as training, leadership, and resource support; and Macroenvironmental (M) — external influences including regulation, economics, and ecosystem community.
2.2.2 Empirical Studies of Signing Usability
@@
The usability of traditional signing tools, often studied in the context of encryption, remains a significant challenge. Whitten and Tygar’s seminal study, “Why Johnny Can’t Encrypt” [99], found PGP 5.0’s interface too complex for non-experts, a problem that persisted in PGP 9 [67], especially around key verification, transparency, and signing. Subsequent work [14, 65, 5] has underscored the public-key model’s inherent complexity, Several works have subsequently examined the use and automation of signing in specific domains, such as social media conversations [28, 64, 8]. Most pertinent to software signing, a recent measurement study by Schorlemmer et al. offers a summative assessment — that traditional software signing tools are unusable [88].
These prior usability studies of signing tools are limited in applicability to software signing. Much of these works have focused on communication contexts such as email and message encryption, where users have less technical expertise [46, 43] than software engineers securing their software supply chains. These studies also examine only traditional signing technologies. We contribute the first usability study of modern “next-generation” software signing tools.
3 Knowledge Gaps & Research Questions
Our literature review shows a gap in formative assessments of software signing tools, Moreover, no studies have examined next-generation signing tools, which are structurally and functionally different from traditional signing tools. We ask:
@@ 3 Knowledge Gaps & Research Questions
    What usability strengths and weaknesses of next-generation software signing tools shape their adoption in practice?
@@
    What factors drive practitioners to adopt, retain, or replace software signing tools over time?
4 Methodology
To investigate our research questions, we adopted a case study approach. Our overall methodology is summarized in Figure˜1. This section details our research design, including the study rationale (section˜4.1), case study context (section˜4.2), instrument design (section˜4.3), data collection (section˜4.4), and analysis procedures (section˜4.5).
@@ 4 Methodology
4.1 Rationale for Case Study
We selected the case study method for our research because it allows for an in-depth exploratory examination of our chosen case, enabling a deeper understanding of the usability factors that influence the adoption of a software signing tool in practice [63, 103, 23, 34], rather than a broad but superficial cross-analysis [60]. In selecting Sigstore, we adhered to the criteria outlined by Yin [103] and Crowe et al. [23]. Specifically, we considered the intrinsic value and uniqueness of the case study unit (high), access to data (adequate), and risks of participation (low). In section˜6 we discuss generalizability.
4.2 Case Study Context – Sigstore
@@
Use in Practice: Sigstore provides keyless signing, where developers authenticate using existing OIDC identities (e.g., GitHub, Google) instead of generating and distributing long-lived keys. It issues short-lived (ephemeral) certificates that simplify key management and reduce long-term security risks compared to traditional signing tools. Within 13 months of its public release in 2021, Sigstore recorded 46 million artifact signatures [36]. It has since gained broad adoption across open-source ecosystems (e.g., PyPI [49], Maven [81], NPM [78]) and major companies (e.g., Autodesk [77], Verizon [76], Yahoo [102]). Given its rapid uptake, integration into critical ecosystems, and pioneering design features, Sigstore provides a compelling case study for examining the usability of next-generation software signing tools.
Components & Workflow: Because our results directly reference specific Sigstore components and workflow steps, we outline Sigstore’s architecture and workflow to orient the reader. Following Newman et al. [104], the components are:
  • Cosign: Command-line and library support (e.g., cosign, sigstore-java) for creating, verifying, and bundling signatures and attestations. Integrations may differ across ecosystems (e.g., GitHub Actions vs. Maven plugins).
  • Identity Provider: Services such as GitHub or Google that authenticate signers and issue short-lived identity tokens (e.g., via OpenID Connect).
@@
  • Monitors and Verifiers: Tools that validate the signer’s identity, certificate status, and signature integrity.
Figure˜2 summarizes the Sigstore software signing workflow. First, the signer authenticates with an OIDC provider to obtain a short-lived token. Next, the certificate authority Fulcio verifies this token and issues a one-time signing certificate. The signer then uses Cosign to apply that certificate to the software artifact. Finally, Rekor logs the certificate and signature, and verifiers retrieve entries from both Fulcio’s log and Rekor to confirm the signer’s identity and ensure the certificate and signature remain valid.
Commonality with other Next-Gen Tools: Sigstore, like other next-generation signing tools such as OpenPubKey [26], AWS Signer [7], and SignServer [73], shares core functionality in signer identity management and automated key handling. However, unlike Sigstore, most of these tools do not include a built-in certificate authority (e.g., OpenPubKey) or a transparency log (e.g., OpenPubKey, SignServer). Instead, such services must be added via external integrations. While our focus on Sigstore allows for an in-depth case analysis, many of its design principles—such as identity-based authentication and ephemeral keys—are shared by its peers, suggesting potential generalizability across next-generation signing tools.
4.3 Instrument Design & Development
Per section˜2.2, our study is a formative usability study. This type of evaluation offers methodological flexibility and prioritizes input from experienced practitioners to better articulate a system’s usability concerns and produce actionable guidance [51]. To elicit long-form, context-rich opinions and detailed observations, we employed semi-structured interviews to examine various aspects of tooling usability (e.g., implementation challenges, perceived benefits, and comparisons to other signing tools), following guidelines from Saldaña [68]. This approach allowed us to pose a consistent core of questions across all participants while probing unique aspects of each participant’s circumstances.
Sigstore is an emerging (though maturing) technology, and case study research recommends exploratory designs in such contexts [103, 63]. To develop our interview protocol, we therefore pursued open, inductive inquiry to capture the full range of practitioner usage experiences. After this exploratory phase, we mapped our findings onto a formal usability framework, ensuring it reflected practitioners’ experiences rather than constraining data collection to preconceived categories (section˜4.5). Therefore, in constructing the protocol, we: